Activity U cannot begin until both activities S and T have been completed. Activity V cannot begin until both activities S ad T have been completed. The duration for U,V, and T are 23,11,70, and 41 days, Respectively. if the earliest start time for S is at the end of day 53 and the earliest start for T is at the end of day 51, the earliest finish time for V is at the end of day______?
Answer #1

Since we have to know earliest finish of V, we will analyse network path for V only. V cannot start till S and T are completed

Early start for S = End of day 53, duration = 70 (Assumed 70 since not clearly mentioned in question)

Early finish of S = 53+70=End of day 123

Early start for T = End of day 51, duration = 41 (Assumed)

Early finish for T = 51+41=End of day 92

Since early finish of S > early finish of T, S governs.

Hence, early finish of V = Early finish of S + Duration of V = 123 + 11 = End of day 134

Answer:134